May I please have some help solving this equation?
Find all solutions of the equation below. (If there are more answer boxes than answers, enter NONE.)
x^4 - 256 = 0

Answer by stanbon(75887)

x^4 - 256 = 0
Factor:
(x^2-16)(x^2+16) = 0
(x-4)(x+4)(x+4i)(x-4i)=0
x = 4 or -4 or -4i or 4i
===============
Cheers,
Stan H.